# Travel Experience in Public Transport Dataset

### v1

@@ Methods
We recommend to use the interactive map to get an overview of the data. We also recommend to use Heart Rate and HRV only filtered by physical activity through transportation mode, with one suggestion being the filtered heart rate we used in the exemplary analyses. To more reliably know whether a participant was physically active we recommend the GPS data, the self-reported mode and the transportation mode detected by the Movinglab app. Furthermore, note that GPS data has a mean accuracy of 8.60m. Not all Experience Sampling data (variables ending with _xs) have a corresponding Longitude and Latitude value, so that filtering by valid Longitude and Latitude values excludes some of the freely reported experience sampling values. This is, therefore, the case for all dataframes but the original ’ExperienceMapsDataset.csv’ when running the code. Both the subjective travel mode report (ModeButton) and the automated travel mode recognition (Mode_keepmovingLast) should be treated with caution, as the respective GPS location indicates that both are not reliable (i.e., participants forgot to indicate their mode change and the Movinglab app sometimes recognized wrong travel modes). The data have limited location variation with the included German cities, and the results’ generalizability needs to be evaluated in future studies.

### v2

@@ 0.0.1 Heart Rate Measures and Permutation Analysis of Heart Rates in Cold and Hot Spots
The accompanying R code reproduces all the results presented in this paper but is not necessary for further use of the dataset. The R-Markdown ’01_Basics-read-translate.Rmd’ and ’02_Preprocessing.Rmd’ contain all the preprocessing of the data, ’03_hotcold spots.Rmd’ contains the code for finding the hot and cold spots of stress and travel satisfaction, and ’04_Rshinyapp.Rmd’ contains the downsampling code and interactive visualization.
Preprocessed cardiac activity data is already available within the provided columns. We recommend to use the interactive map to get an overview of the data. We also recommend to use Heart Rate and HRV only filtered by physical activity through transportation mode. For a more reliable determination of participants’ physical activity, we suggest integrating and interpreting GPS data, self-reported transportation modes, and the modes detected by the MovingLab app. All original cardiac activity data are included in the data. The MAD-transformed values of RMSSD and heart rate are one suggestion to exclude cardiac activity data during and one minute after physically active traveling modes. Literature suggests that a sitting period of 1 to 10 minutes is sufficient for analyzing cardiac activity data after physical activity [50, 54]. The specific duration within this range should be determined by the researcher, depending on the desired level of strictness in discarding potentially affected data. As the times of physical activity last usually at least a few minutes, we do not recommend to use any imputations.
Note that not all Experience Sampling data have a corresponding Longitude and Latitude value, so that filtering by valid Longitude and Latitude values excludes some of the freely reported experience sampling values. This is, therefore, the case for all dataframes but the original ’ExperienceMapsDataset.csv’. For translation of the columns with free text input we recommend to use https://www.deepl.com/de/translator. As demonstrated in the technical validation of travel mode detection, none of the three travel mode columns - ModeButton (subjective reporting), Mode_keepmoving (from the DLR app), and ActivityClass (from the chest belt) — achieved perfect accuracy. Errors occurred due to participants forgetting to indicate mode changes and instances where the app or chest belt incorrectly recognized travel modes. To address these limitations, we recommend leveraging a combination of all three detection methods. For instance, subjective mode reports can be visualized on a map and cross-validated against the plausibility of travel modes based on street types. If inconsistencies are detected, the DLR app and ECG chest belt travel mode detections can provide additional context. Furthermore, for imputing missing GPS data, the open-source routing software Valhalla offers a viable solution (https://github.com/valhalla).
To include more context data an option is to use OpenStreetMaps’ export data which contains raw geographic and metadata information about nodes, ways, and relations. To obtain weather information, we recommend to use One Call API 3.3 from OpenweatherMap (https://openweathermap.org/api/one-call-3). As Figure 5 indicates, satisfaction levels seem to be generally lower in November than the rest of the year, and a correlation between weather and satisfaction data could be interesting.
The dataset is geographically limited, including data only from the German cities of Braunschweig and Wolfenbüttel. Consequently, the generaliezability of the findings must be further evaluated in future studies employing similar data acquisition methods. To enhance the robustness of future data collection, the approach could be expanded to include survey items inspired by datasets such as the Household, Income, and Labor Dynamics in Australia (HILDA) dataset [55, 56]. While efforts were made to balance the sample by gender, participants were drawn exclusively from our institution’s participant pool, which may skew the demographics toward individuals with prior experience in research studies or interested in the research topics as well as those motivated by financial incentives. Additionally, the sample’s age distribution and educational or occupational diversity were not controlled, which could introduce biases in interpreting results (see Tables 2, 3 for demographic data). Self-selection bias of participants is more likely in experience sampling studies than other studies due to the intense time-demands on participants - and certain types of individuals might answer more frequently than others, even though incentivization by money has shown to be a good solution to this [26]. The relatively small sample size (44 participants) represents a limitation in terms of the generalizability of the findings. The insights from this dataset should therefore be interpreted with caution, particularly when considering its applicability to broader populations. However, the sample size aligns with the exploratory nature of the study, which focuses on integrating subjective, cardiac activity, and contextual data to investigate travel experiences. Future studies should aim to replicate these methods with larger and more representative samples to validate and extend these findings.